# Kibble-Zurek mechanism and errors of gapped quantum phases

### v1

@@ I Introduction
Recent advancements in the development of several quantum architectures has propelled significant interest in exploring various quantum many-body phenomena Smith et al. (2019); de Léséleuc et al. (2019); Brydges et al. (2019); Bhattacharyya et al. (2022); Wei et al. (2023). In particular, QKZM has been validated on different quantum hardware platforms by estimating the critical properties, especially in Ising-like models Keesling et al. (2019); Dupont and Moore (2022). Motivated by the recent progress, in this work, we introduce a novel numerical scheme to obtain the scaling exponent in QPTs that are in principle accessible on a quantum device. We emphasize that the introduced method is not limited to the scope of the Landau symmetry breaking principle, and can be applied to a broader range of QPTs involving phases characterized by non-local order. To this extent, we further present strategies that enable numerical and experimental observation of QKZM involving symmetry protected topological phases. This in turn can be used as a scheme to profile and benchmark the performance of quantum computing devices.
@@ II.1 Revisiting the QKZM
The QKZM process relies on a quench dynamics that can be generated by a time-dependent Hamiltonian, $H(t)$, connecting point $I$ to $F$,
@@
with $g(t)=-t/\tau_{Q}$ where $t\in(-\infty,0]$, $\tau_{Q}$ being the rate of the quench del Campo and Zurek (2014). Having the system prepared in the groundstate of the Hamiltonian, $H_{I}$, we evolve the system under the total Hamiltonian mentioned in Eq. 1. Assuming that there exists a second-order QPT at some critical strength $g(t_{c})$, QKZM establishes a power law relation between the defect density, $\eta$, at the final time and the quench rate, $\tau_{Q}$. The resulting relation is expressed as $\eta\propto\tau_{Q}^{-\mu}$, where $\mu$ is the scaling exponent characterizing the universality of the underlying QPT  Zurek et al. (2005); Dziarmaga (2005b); Polkovnikov (2005).
@@ II.2 Methods to compute defect density
Figure 1: QKZM with the quench protocol driven across a paramagnetic-ferromagnetic QPT as in the tranverse field Ising model. Defect density, $\eta$, as a function of the quench rate, $\tau_{Q}$, with $\eta$ computed using the strategy based on (a) expectation value of the projector, (b) Monte-Carlo sampling. The linear scaling remains indistinguishable for progressively increasing values of $N$ indicating convergence to the behavior in the thermodynamic limit. We note that in the rest of the figures, we consider the y-axis on a decimal logarithm scale for better representation but the actual fit is made with respect to $\log_{2}$ on both axes. (Inset) By performing finite-size scaling analysis we estimate the critical exponent $\mu$ to be (a) 0.502(1), (b) 0.504(2) in the thermodynamic limit.
@@ III Quantum phases with local order
The above Hamiltonian exhibits a QPT with the groundstate being a paramagnet in the low perturbed regime while being a ferromagnet in the high perturbed regimes with a criticality at $g(-\tau_{Q})$=1. To estimate the critical exponent, we first choose the initial state to be the groundstate of Eq. 4 at some high field strength, thereby belonging to the paramagnetic phase. Then we evolve the system across the QPT guided by the quench protocol. In the current context, the reference state that leads to the construction of errors is the ferromagnetic groundstate i.e., |$\ket{0000...00}$ and |$\ket{1111...11}$. The presence of the transverse field gives rise to errors that are recognized by neighboring spins with opposite parity when measured in the $\sigma_{z}$ basis. Having introduced the quench dynamics and the errors associated with the ferromagnetic groundstate, we quantify the scaling exponent by employing the strategies as outlined in Sec. II.2.
@@ IV.1 Su–Schrieffer–Heeger (SSH) model
with $v(t)=-t/\tau_{Q}$, $t\in(-\infty,0]$ with $\tau_{Q}$ being the quench rate. The time independent version of Eq. 6 with $v(t)=v$ is exactly solvable in the case of periodic boundary conditions. It hosts gapped phases in the extremal limits of $v\ll w$ and $v\gg w$ with the gap closing at $v=w$. In the case of open boundaries, in the limit of $v\ll w$ it is known that the topological phase is identified by the presence of edge modes, characterized as non-trivial SPT phase. While in the other limit the phase remains topological with no edge modes, characterized as trivial SPT phase with the phase transition occurring at $v=w$ Elben et al. (2020); Jamadagni and Weimer (2022b). For the rest of the analysis we consider the SSH Hamiltonian on a 1D-chain i.e., with open boundaries.
@@ IV.2 Extended SSH model
Figure 4: (a) Qualitative groundstate phase diagram of the extended SSH model with arrows indicating the quench dynamics through different QPTs (b, c, d) from trivial to SPT phase by fixing $\delta/w$ to be (b) 0.5, (c) 1.0, (d) 3.0, and (e) from antiferromagnetic (AFM) to SPT phase by fixing $w/v=2$. In the former, the linear scaling of defects (on a $\log-\log$ scale) is in accordance with the QKZM (inset) resulting in the critical exponent, $\mu$ to be (b) 0.4767(1), (c) 0.4328(1), (d)0.521(1). (e) The finite size effects in the AFM-SPT quench are more predominant and are reflected in the deviations from the linear scaling included in the error-estimation of (inset) critical exponent, $\mu$=0.50(2).
@@ IV.2.1 Trivial to SPT transition
We dynamically traverse across the trivial to non-trivial SPT topological transition by employing a quench protocol i.e., by setting $v$ to be $v(t)=-t/\tau_{Q}$, $t\in(-\infty,0]$ and $\delta$ to be a constant belonging to the set $\{0.5,1,3\}$ in Eq. 9. The initial state is chosen as the groundstate of the Hamiltonian in the regime of $v\gg w$ with the final evolved state belonging to the non-trivial SPT phase. The reference state is as introduced in the earlier section via Eq. 7. The above setting leads us to a final evolved state with total defect density given by $\eta$, as in Eq. 8. In Fig. 4, we estimate the critical exponents corresponding to $\delta$ in $\{0.5,1,3\}$. Our numerical experiments show that with interactions turned on, the value of the scaling exponent deviates significantly from 0.5 as in the non-interacting case. We also notice that the value decrease along the transition line $w/v$=1 upto the point where the three phases co-exist and then recovers to 0.5 beyond that.
@@ IV.2.2 AFM to SPT transition
The phase diagram allows for the exploration of critical exponent associated with the AFM-SPT phase transition. To this end, the quench protocol is defined by $\delta(t)=-t/\tau_{Q}$ with $t\in(-\infty,0]$ while fixing $w/v=2$ in Eq. 9. The initial state for evolution is chosen to be the groundstate of $H_{\text{eSSH}}$ in the limits of $\delta\gg w/v$, thereby being smoothly connected to an antiferromagnet. As the system is driven into a non-trivial SPT phase, the reference state remains the same. It might be intuitive to conclude that the total error density is given by Eq. 8. However, on the contrary this is not the case as there are additional corrections involved. It is important to note that the final evolved state encapsulates the errors of the groundstate at finite $w/v$ that need to be subtracted from Eq. 8 to obtain the accurate defect densities. In Fig. 4, we note that by incorporating the additional correction terms we substantiate the predictions as in the QKZM and further obtain the scaling exponent.
@@ IV.3 Cluster state model
where $g(t)$ is as defined in Eq. 1. For the rest of the analysis, we consider the Hamiltonian on a 1D spin chain i.e., with open boundary conditions. In the case of time-independent perturbation, the Hamiltonian hosts a SPT phase at low perturbation strength and a paramagnetic phase at high perturbation strength with a QPT at some perturbation strength Cong et al. (2019). In the absence of perturbation, the groundstate of the Hamiltonian is short-range entangled and protected by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry Son et al. (2011). The groundstate is also referred to as 1D cluster state with applications in measurement-based quantum computing Nielsen (2006); Deng et al. (2017).
@@
Our numerical analysis shows the critical exponent to be $\mu=0.493(2)$, see Fig. 5. This establishes that the numerical methods introduced here work for wider class of gapped phases that demand multi-site interactions.
@@ V Quantifying criticality using digital quantum computers
Figure 6: Simulating the quench dynamics involving a topological phase transition as in the extended SSH model on a digital quantum computer. (a) The TEBD protocol comprising of the two-qubit unitary, $U$, see Eq. 12. (b) The decomposition of $U(\alpha,\beta,\gamma)$ in terms of single qubit rotations where $\theta=\frac{\pi}{2}-2\gamma$, $\phi=2\alpha-\frac{\pi}{2}$, $\lambda=\frac{\pi}{2}-2\beta$ and CNOTs. The predicted power law scaling in the QKZM is recovered by considering (c) partial defect density i.e., density fluctuations estimated by sampling in $\sigma_{z}$ basis similar to shot data in experiments, (d) partial defect density by computing the expectation value of the density fluctuations, (e) total defect density determined using the projector expectation value. (Inset) The scaling exponent, $\mu$ obtained by performing finite size scaling resulting in (c) 0.509(3), (d) 0.507(1), (e) 0.508(1) that are in good agreement among themselves while also to the quench protocol employed in Sec. IV.
@@
The above can be realized on a digital quantum architecture by decomposing it further into single-qubit Pauli rotations and two-qubit Controlled-NOT (CNOT) gates Vatan and Williams (2004); Smith et al. (2019), see Fig. 6(b). We note that the procedure outlined in this section and the simulations thereof, assume an ideal quantum computing architecture. Relaxing the above constraints involves employing additional techniques for instance, fine-tuned finite-size scaling analysis, employing circuit depth reduction techniques to achieve the required evolution in shorter time, and integrating error mitigation strategies, to mention a few. We leave this exploration in the context of topological phases to the future.
@@ V.2 Computing partial and total defect densities
Figure 7: Monte-Carlo sampling to obtain defect densities estimated by considering (a) all errors, (b) density errors, (c) phase errors for the extended SSH model driven into the non-trivial SPT phase. (Inset) Estimating the corresponding critical exponents by performing finite-size analysis leading to $\mu=$ (a) 0.433(1), (b) 0.434(1), (c) 0.435(2) that exhibit agreement with each other and also with the critical exponent obtained based on projector expectation values as in Sec. IV.
@@ Acknowledgements.
AJ would like to thank Andreas Läuchli for fruitful discussions. AB would like to thank the FISPAC Research Group, Department of Physics, University of Murcia, especially, Jose J. Fernández-Melgarejo, for hospitality during the course of this work. AB is supported by the Mathematical Research Impact Centric Support Grant (MTR/2021/000490) by the Department of Science and Technology Science and Engineering Research Board (India) and the Relevant Research Project grant (202011BRE03RP06633-BRNS) by the Board Of Research In Nuclear Sciences(BRNS), Department of Atomic Energy (DAE), India. AB also acknowledge associateship program of Indian Academy of Science, Bengaluru.

### v2

@@ I Introduction
Recent advancements in the development of several quantum architectures has propelled significant interest in exploring various quantum many-body phenomena Smith et al. (2019); de Léséleuc et al. (2019); Brydges et al. (2019); Bhattacharyya et al. (2022); Wei et al. (2023). In particular, QKZM has been validated on different quantum hardware platforms by estimating the critical properties, especially in Ising-like models Keesling et al. (2019); King et al. (2021); Dupont and Moore (2022); King et al. (2022); Bando et al. (2020); Du et al. (2023). Motivated by the recent progress, in this work, we introduce a novel numerical scheme to obtain the scaling exponent in QPTs that are in principle accessible on a quantum device. We emphasize that the introduced method is not limited to the scope of the Landau symmetry breaking principle, and can be applied to a broader range of QPTs involving phases characterized by non-local order. To this extent, we further present strategies that enable numerical and experimental observation of QKZM involving symmetry protected topological phases. This in turn can be used as a scheme to profile and benchmark the performance of quantum computing devices.
@@ II.1 Revisiting the QKZM
The QKZM provides a theoretical framework for understanding the behavior of a system undergoing a continuous QPT when it is driven out of equilibrium by a rapid change in a control parameter. The mechanism relies on a quench dynamics that can be generated by a time-dependent Hamiltonian, $H(t)$, connecting point $I$ to $F$,
@@
with $g(t)=-t/\tau_{Q}$ denotes a linear quench with the rate $\tau_{Q}$ in the time interval $t\in(-\infty,0]$ del Campo and Zurek (2014). Having the system prepared in the groundstate of the Hamiltonian, $H_{I}$, we evolve the system under the total Hamiltonian mentioned in Eq. 1. Assuming that there exists a second-order QPT at some critical strength $g(t_{c})$, the correlation length $\xi$ diverges at criticality as $\xi\propto|g(t)|^{-\nu}$ with $\nu$ being the correlation-length critical exponent. This is associated with the closing of the energy gap, or the divergence of the relaxation time as $\tau\simeq\Delta E^{-1}\propto|g(t)|^{-z\nu}$ where $z$ is the dynamical critical exponent, see for example Ref. Sachdev, 2011; Dziarmaga, 2005b on procedures to estimate the critical exponents. The QKZM establishes a relation between the the size of the non-adiabatic domains and the critical exponents expressed as
$\xi\propto\tau_{Q}^{\frac{z}{1+z\nu}}.$ (2)
Consequently the defect density in a $D$ dimensional system exhibits a power-law scaling as $\eta\propto\tau_{Q}^{-\mu}$ where $\mu=Dz/(1+z\nu)$ is the QKZM scaling exponent determined by the universality class of the underlying QPT  Zurek et al. (2005); Dziarmaga (2005b); Polkovnikov (2005). In this work, we especially focus on one-dimensional systems, however our method can be extended to generic gapped phases, see App. H where we sketch the outline for two dimensional systems.
@@ II.2 Methods to compute defect density
(a) Refer to caption (b) Refer to caption
Figure 1: QKZM with the quench protocol driven across a paramagnetic-ferromagnetic QPT as in the tranverse field Ising model. Defect density, $\eta$, as a function of the quench rate, $\tau_{Q}$, with $\eta$ computed using the strategy based on (a) expectation value of the projector, (b) Monte-Carlo sampling. The linear scaling remains indistinguishable for progressively increasing values of $N$ indicating convergence to the behavior in the thermodynamic limit. We note that in the rest of the figures, we consider the y-axis on a decimal logarithm scale for better representation but the actual fit is made with respect to $\log_{2}$ on both axes. (Inset) By performing finite-size scaling analysis we estimate the critical exponent $\mu$ to be (a) 0.502(1), and (b) 0.504(2) in the thermodynamic limit.
@@ III Quantum phases with local order
The above Hamiltonian exhibits a QPT with the groundstate being a paramagnet in the low perturbed regime while being a ferromagnet in the high perturbed regimes with a criticality at $g(-\tau_{Q})$=1. The paramagnet-ferromagnet transition belongs to the Ising universality class with the critical exponents $\mu=z=1.0$ Sachdev (2011). To estimate the critical exponent, we first choose the initial state to be the groundstate of Eq. 5 at some high field strength, thereby belonging to the paramagnetic phase. Then we evolve the system across the QPT guided by the quench protocol. In the current context, the reference state that leads to the construction of errors is the ferromagnetic groundstate i.e., |$\ket{0000...00}$ and |$\ket{1111...11}$. The presence of the transverse field gives rise to errors that are recognized by neighboring spins with opposite parity when measured in the $\sigma_{z}$ basis. Having introduced the quench dynamics and the errors associated with the ferromagnetic groundstate, we quantify the scaling exponent by employing the strategies as outlined in Sec. II.2.
@@ IV.1 Su–Schrieffer–Heeger (SSH) model
with $v(t)/w=-t/\tau_{Q}$, $t\in(-\infty,0]$ with $\tau_{Q}$ being the quench rate. The time independent version of Eq. 7 with $v(t)=v$ is exactly solvable in the case of periodic boundary conditions. It hosts gapped phases in the extremal limits of $v\ll w$ and $v\gg w$ with the gap closing at $v=w$. In the case of open boundaries, in the limit of $v\ll w$ it is known that the topological phase is identified by the presence of edge modes, characterized as non-trivial SPT phase. While in the other limit the phase remains topological with no edge modes, characterized as trivial SPT phase with the phase transition occurring at $v=w$ Elben et al. (2020); Jamadagni and Weimer (2022b). For the rest of the analysis we consider the SSH Hamiltonian on a 1D-chain i.e., with open boundaries.
@@ IV.2 Extended SSH model
Figure 4: (a) Qualitative groundstate phase diagram of the extended SSH model with arrows indicating the quench dynamics through different QPTs (b, c, d) from trivial to SPT phase by fixing $\delta/w$ to be (b) 0.5, (c) 1.0, (d) 3.0, and (e) from antiferromagnetic (AFM) to SPT phase by fixing $w/v=2$. In the former, the linear scaling of defects (on a $\log-\log$ scale) is in accordance with the QKZM (inset) resulting in the critical exponent, $\mu$ to be (b) 0.4767(1), (c) 0.4328(1), and (d)0.521(1). (e) The finite size effects in the AFM-SPT quench are more predominant and are reflected in the deviations from the linear scaling included in the error-estimation of (inset) critical exponent, $\mu$=0.50(2).
@@ IV.2.1 Trivial to SPT transition
We dynamically traverse across the trivial to non-trivial SPT topological transition by employing a quench protocol i.e., by setting $v$ to be $v(t)=-t/\tau_{Q}$, $t\in(-\infty,0]$ and $\delta$ to be a constant belonging to the set $\{0.5,1,3\}$ in Eq. 10. The initial state is chosen as the groundstate of the Hamiltonian in the regime of $v\gg w$ with the final evolved state belonging to the non-trivial SPT phase. The reference state is as introduced in the earlier section via Eq. 8. The above setting leads us to a final evolved state with total defect density given by $\eta$, as in Eq. 9. In Fig. 4, we estimate the critical exponents corresponding to $\delta$ in $\{0.5,1,3\}$. Our numerical experiments show that with interactions turned on, the value of the scaling exponent deviates significantly from 0.5 as in the non-interacting case. We also notice that the value decrease along the transition line $w/v$=1 upto the point where the three phases co-exist and then recovers to 0.5 beyond that. To the best of our knowledge, our studies is one of the first to report such estimates.
@@ IV.2.2 AFM to SPT transition
The phase diagram allows for the exploration of critical exponent associated with the AFM-SPT phase transition. To this end, the quench protocol is defined by $\delta(t)=-t/\tau_{Q}$ with $t\in(-\infty,0]$ while fixing $w/v=2$ in Eq. 10. The initial state for evolution is chosen to be the groundstate of $H_{\text{eSSH}}$ in the limits of $\delta\gg w/v$, thereby being smoothly connected to an antiferromagnet. As the system is driven into a non-trivial SPT phase, the reference state remains the same. It might be intuitive to conclude that the total error density is given by Eq. 9. However, on the contrary this is not the case as there are additional corrections involved. It is important to note that the final evolved state encapsulates the errors of the groundstate at finite $w/v$ that need to be subtracted from Eq. 9 to obtain the accurate defect densities. To be more precise, let us denote the final evolved state by |$\ket{\psi}_{f}$ and the ground state at $w/v$ = 2 by |$\ket{\psi}_{g}$, with the corresponding defect density given by $\eta_{f}$ and $\eta_{g}$ respectively. The true defect density involving the additional corrections terms is therefore given by the following
[…]
It is clear to see that $\eta_{\text{eff}}$ is non-negative as the overlap of the |$\ket{\psi_{g}}$ with the reference state $|\minusket\minus\ket{\boldsymbol{\minus}}$ is more in comparison to the final evolved state, |$\ket{\psi_{f}}$ that involves additional excitations. In Fig. 4, we note that by incorporating the additional correction terms we substantiate the predictions as in the QKZM and further obtain the scaling exponent.
@@ IV.3 Cluster state model
where $g(t)$ is as defined in Eq. 1. For the rest of the analysis, we consider the Hamiltonian on a 1D spin chain i.e., with open boundary conditions. In the case of time-independent perturbation, the Hamiltonian hosts a SPT phase at low perturbation strength and a paramagnetic phase at high perturbation strength with a QPT at some perturbation strength Cong et al. (2019). In the absence of perturbation, the groundstate of the Hamiltonian is short-range entangled and protected by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry Son et al. (2011a). The groundstate is also referred to as 1D cluster state with applications in measurement-based quantum computing Nielsen (2006); Deng et al. (2017).
@@
Our numerical analysis shows the critical exponent to be $\mu=0.493(2)$, see Fig. 5. This establishes that the numerical methods introduced here work for wider class of gapped phases that demand multi-site interactions. Estimates for the critical exponents are obtained for different perturbed model involving the cluster state Hamiltonian leading to the same scaling exponent as in our case Son et al. (2011b).
@@ V Quantifying criticality using digital quantum computers
Figure 6: Simulating the quench dynamics involving a topological phase transition as in the extended SSH model on a digital quantum computer. (a) The TEBD protocol comprising of the two-qubit unitary, $U$, see Eq. 14. The decomposition of $U(\alpha,\beta,\gamma)$ in terms of single qubit rotations and CNOTs is as in Refs. Vatan and Williams, 2004; Smith et al., 2019. The predicted power law scaling in the QKZM is recovered by considering (b) partial defect density i.e., density fluctuations estimated by sampling in $\sigma_{z}$ basis similar to shot data in experiments, (c) partial defect density by computing the expectation value of the density fluctuations, and (d) total defect density determined using the projector expectation value. (Inset) The scaling exponent, $\mu$ obtained by performing finite size scaling resulting in (b) 0.504(2), (c) 0.505(1), and (d) 0.507(1) that are in good agreement among themselves while also to the quench protocol employed in Sec. IV
@@
The above can be realized on a digital quantum architecture by decomposing it further into single-qubit Pauli rotations and two-qubit Controlled-NOT (CNOT) gates Vatan and Williams (2004); Smith et al. (2019), We note that the procedure outlined in this section and the simulations thereof, assume an ideal quantum computing architecture. Relaxing the above constraints involves employing additional techniques for instance, fine-tuned finite-size scaling analysis, employing circuit depth reduction techniques to achieve the required evolution in shorter time, and integrating error mitigation strategies, to mention a few. We leave this exploration in the context of topological phases to the future.
@@
Figure 7: Estimating the scaling exponent by introducing SWAP noise. We recover the power law scaling and the associated exponent in the case of half-swap noise by computing the defect density using projector expectation value of (a) all errors and (b) density fluctuations resulting in $\mu=$ (a) 0.5081(1) (b) 0.5037(5) respectively. Similarly, in the context of full-swap noise, the predictions of QKZM remains valid as we recover the scaling exponent as in the noiseless evolution by computing the defect density using projector value of (c) all errors and (d) density fluctuations resulting in $\mu=$ (c) 0.5075(3), (d) 0.5023(8) respectively.
@@ V.2 Computing partial and total defect densities
V.3 Estimating critical exponent under error invariant noisy evolution
In this section, we show that access to the errors of gapped phases allow the exploration of noisy dynamics (in this case coherent noise) that respect the QKZM. In the current scenario, we note that the gapped errors given by {|$\plus\{\ket{\boldsymbol{0}},\ket{\boldsymbol{1}},\ket{\boldsymbol{\plus}}\}$ remain invariant under the action of a SWAP gate. This further leads us to the notion that dynamics of the extended SSH model driven across a trivial non-trivial topological phase interspersed with SWAP unitaries leave the scaling exponent invariant as in the noiseless scenario thereby recovering the prediction of the QKZM. To validate the above, we consider the evolution as in the TEBD protocol and inject SWAPs unitaries leading to noisy evolution. We study two different SWAP injection protocols and label them as half-swap and full-swap injection. To motivate the above protocols we consider the TEBD algorithm as in Fig. 6(a) with a half (full)-swap injection implying a SWAP is prepended to all the unitaries in block identified by 2 (1 and 2 blocks). In both the scenarios, we recover the scaling exponent as in the noiseless evolution obtained by computing the expectation value as in Eq. 9, see Fig. 7(a), (c). In addition, we also note that this can be achieved on a digital quantum computer as the density fluctuations i.e., shot measurement in the $\sigma_{z}$ basis, also recover the expected scaling exponent, see Fig. 7(c), (f). In summary, we observe that unitaries that leave the error space invariant, when injected into the quench dynamics retain the prediction of the QKZM thereby leading to same scaling exponent as in the absence of unitary injection.
@@ Acknowledgements.
AJ would like to thank Andreas Läuchli for fruitful discussions. AB would like to thank the FISPAC Research Group, Department of Physics, University of Murcia, especially, Jose J. Fernández-Melgarejo, for hospitality during the course of this work. AB is supported by the Mathematical Research Impact Centric Support Grant (MTR/2021/000490) by the Department of Science and Technology Science and Engineering Research Board (India) and the Relevant Research Project grant (202011BRE03RP06633-BRNS) by the Board Of Research In Nuclear Sciences(BRNS), Department of Atomic Energy (DAE), India. AB also acknowledge associateship program of Indian Academy of Science, Bengaluru. We have also uploaded sample code snippets at the following repository that enables the reproduction of the results: https://github.com/amitjamadagni/KZ_qdyn.